OpenStudy (anonymous):

For a class pizza party, Esther bought 9 bottles of juice and twelve bottles of water. Let j represent the cost of juice and w represent the cost of water. Write an expression to determine the total cost for the drinks. Then, calculate the cost if juice costs two dollars and water costs one dollar.

OpenStudy (anonymous):

C = 9j + 12w C = 9(2) + 12(1) = 30

OpenStudy (anonymous):

nine j plus twelve w; thirty-three dollars twelve j plus nine w; thirty-three dollars nine j plus twelve w; thirty dollars twelve j plus nine w; thirty dollars

OpenStudy (anonymous):

nine j plus twelve w; thirty dollars twelve j plus nine w; thirty dollars either one looks correct.